COMPUTATIONAL MOLECULAR SPECTROSCOPY OF FeCO IN THE X̃^{3}Σ^{−} AND 1^{5}Σ^{−} ELECTRONIC STATES

\underline{TSUNEO HIRANO}, REI OKUDA, and UMPEI NAGASHIMA, Research Institute for Computational Sciences, National Institute of Advanced Industrial Science and Technology, 1-1-1 Umezono, Tsukuba, Ibaraki 305-8568, Japan; PER JENSEN, Theoretische Chemie, Bergische Universität, D-42097 Wuppertal, Germany.

FeCO is a molecule of astrophysical interest. It also served as a bench mark molecule for testing basis sets and ab initio calculation methods.^{a}^{,}^{b} We have previously reported molecular constants and relative energies of X̃^{3}Σ^{−} and 1^{5}Σ^{−} states of FeCO, based on the two-dimensional ab initio potential energy surfaces (PESs) determined at the MR-SDCI+Q+E_{rel}/[Roos ANO (Fe, C, O)] and MR-ACPF+E_{rel}/[Roos ANO (Fe, C, O)] levels of theory.^{c} We will report here molecular properties derived from the three-dimensional PESs calculated at the level of MR-SDCI+Q+E_{rel}/[Roos ANO (Fe, C, O)]. Calculated bond lengths r_{e}(Fe–C), r_{e}(C–O), and dipole moment (with experimental r_{s} bond lengths^{d} in parentheses) are 1.722 (1.7270) Å, 1.160 (1.1586) Å, and 3.20 D for the X̃^{3}Σ^{−} state, and 1.844 Å, 1.153 Å, and 0.29 D for the 1^{5}Σ^{−} state, respectively. The relative energy of 1^{5}Σ^{−} has been calculated to be 1.27 kcal mol^{−1}, to be compared with the experimental value of 3.24 kcal mol^{−1}.^{e} The general trends in low-spin/high-spin issue reported for CoH,^{f} CoCN,^{g} and NiCN^{h} are also observed for the NiCN X̃^{3}Σ^{−} and 1^{5}Σ^{−} state pair.

aA. Ricca and C. W. Bauschlicher, Theor. Chem. Acc. 106, 314, (2001).

bT. Noro, M. Sekiya, T. Koga, and H. Matsuyama, Theor Chem. Acc. 104, 146, (2000).

cM. Amano, S.S. Itono, T. Hirano et al., 57th OSU International Symposium on Molecular Spectroscopy, RF11, 2002.

dK. Tanaka, K. Sakaguchi, and T. Tanaka, J. Chem. Phys., 106, 2118 (1997).

eP. W. Villalta and D. G. Leopold, J. Chem. Phys. 98, 7730, (1993).

fM. Tomonari, R. Okuda, U. Nagashima, K. Tanaka, and T. Hirano, J. Chem. Phys., 126, 144307 (2007).

gT. Hirano, Rei Okuda, U. Nagashima, and P. Jensen, J. Chem. Phys., 127, 014303 (2007).

hT. Hirano, Rei Okuda, U. Nagashima, and P. Jensen, Chem. Phys., (2008), in press.